# Hamiltonian learning quantum magnets with non-local impurity tomography

### v1

@@ I Introduction
Quantum magnetism provides a highly flexible playground to engineer exotic phenomena [1, 2]. While a variety of natural materials host frustrated and magnetic states, accessing the most unconventional regimes often requires using artificial platform where Hamiltonians can be engineered [3]. Atomic-scale quantum magnets [4, 5, 6, 7, 8, 9, 10, 11] have opened new avenues for exploring fundamental quantum phenomena, enabling controllable platform to engineer symmetry breaking [12], quantum criticality [13], emergent excitations [14, 15], and topological states [16, 17]. Scanning tunneling microscopy (STM) enables creating and probing artificial quantum states with high spatial and energy resolution, providing detailed insights into the behavior of quantum magnets [4, 5, 6, 7, 8, 18, 19, 9, 10, 11]. Electrically driven paramagnetic resonance with STM have brought quantum control sensing at the atomic scale to remarkable levels. ESR-STM enabling energy resolution several orders of magnitude below thermal noise [10, 20], allowed quantum gate at the atomic level [21], and probing time evolution of magnetic dynamics with atomic accuracy [22]. Single atom manipulation in quantum magnet thus enables a degree of control radically different than in bulk materials, motivating the development of new strategies to tackle open problems in quantum magnetism [23].
@@
Understanding an experimental quantum many-body systems relies on having an accurate description of the many-body interactions. Traditional methods of determining Hamiltonian parameters often rely on fitting experimental data to theoretical models. Characterizing quantum spin liquids represents a remarkable problem in quantum materials [1, 2] due to the intricate nature of its ground state, and often requiring new strategies to find their experimental signatures [24, 25, 26, 27, 28, 29, 30]. As a result, even in relatively simple quantum magnets, determining the Hamiltonian from experimental data can be remarkably difficult [3]. Machine learning provides a flexible strategy to enable learning of Hamiltonian parameters directly from physical observables [23, 31, 32, 33, 34]. Hamiltonian learning can be done with data generated from theoretical models, that can be later on directly applied to experimental measurements [35, 36]. This strategy leverages the exceptional ability of machine learning methods to identify non-trivial correlations in observables that enable extracting information about the underlying ground state and Hamiltonian [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
@@
Here we show how the Hamiltonian of a quantum magnet can be directly inferred from measurements of the response of the spin chain to individual impurities. Our strategy relies on using the response of the system to extract the Hamiltonian using a machine learning strategy. The inclusion of individual impurities create a perturbation in the ground state that directly reflects the underlying Hamiltonian, a correlation that can be mapped with a supervised neural network. We show that this methodology enables extracting long range and anisotropic exchange interactions, often a challenge to infer in experiments. We show that this methodology can be understood as a generalization of Hamiltonian extraction in a fermionic model, a procedure that it is often performed in STM via quasiparticle interference. Our manuscript is organized as follows. In Section II we describe the experimental strategy to obtain the data required for our Hamiltonian extraction. Section III details the many-body spin models we study, including the fermionic version that provides a minimal exemplification of our methodology. The methods employed for Hamiltonian parameter extraction using neural networks are described in Section IV, where we discuss the training process, data generation, and the inclusion of noise to simulate experimental conditions. In Section V, we present the results of the Hamiltonian inference, demonstrating the resilience of our methodology to noise both in the evaluation as well as during training. Finally, in Section VI, we summarize our conclusions.
@@ II Impurity perturbations and non-local response
Our methodology will rely on using the non-local response of the system to extract the Hamiltonian. However, STM measures local quantities rather than non-local ones, and thus first an strategy must be developed to extract the required non-local information. Non-local correlators of quantum spin chain determine the response of the quantum many-body system to perturbation in the sites, and thus are reflected when adding additional impurities that act as exchange sources. Specifically, additional sites featuring strong single ion anisotropy as local perturbations, which then enable to measure elsewhere in the system what is the change they created to the ground state. Experimentally, this was demonstrated with dysprosium single atom magnets [49], enabling atomic-scale magnetic field with stabilities above days. The Dy impurity creates a local exchange field in the neighboring site, which as a result modified the expectation value of the magnetization in the rest the system. This strategy is exemplified in Fig. 1a.
@@
$H_{0}=\sum_{n}J_{1}(1+\delta(-1)^{n})\boldsymbol{S}_{n}\cdot\boldsymbol{S}_{n}+\sum_{k=2}^{5}\sum_{n}J_{k}\boldsymbol{S}_{n}\cdot\boldsymbol{S}_{n+k}$ (7)
, which features a many-body gap for $\delta\neq 0$ and $J_{k}\ll J_{1}$ for $k=2,3,4,5$. Figure 2a shows the magnetization in the different sites upon adding the local exchange perturbation, whereas Figure 2b shows the non-local spin-spin correlator in the absence of perturbation. It is clearly observed that both quantities feature very good correspondance over several orders of magnitude, which allows to infer non-local correlators from measuring local magnetizations of perturbed quantum magnet. The system features a decay corraltion and magnetization for all the different couplings considered, yet interestingly the type of long range coupling included dramatically changes the speed of the decay and the oscillations. In this fashion, even though all those system feature similar many-body gaps, the magnetization and correlators provide a strong fingerprint of the underlying Hamiltonian. These unique behavior for the different coupling parameters is what provides the starting ground to attempt Hamiltonian learning from the observables in Figure 2.
The previous methodology exemplifies how non-local spin correlations can be extracted from local STM measurements by engineering local magnetic sites [4, 5, 6, 54]. As a result, Hamiltonian learning could be performed either directly by measuring the local response a system when adding local perturbations on different sites (Fig. 2a), or with the non-local correlator in the unperturbed limit (Fig. 2b). It is worth noting that this strategy can be also used for single particle systems, just by replacing the spin operators by fermionic operators and adapting the Kubo formalism accordingly. In terms of Hamiltonian learning, we will illustrate both single particle and the quantum many-body magnet cases. For the sake of concreteness, we will perform Hamiltonian learning with local measurements of purbed systems for the fermionic case, and with the non-local correlators for the quantum many-body limit, yet noting that either local perturbed measurements or non-local unperturbed measurements allow for Hamiltonian learning in both cases. It is finally worth noting that time-dependent measurements, either local or non-local, could be used for Hamiltonian learning[55, 56, 57, 58, 59, 60, 61, 62, 63, 64].
@@ III Models
In this section, we introduce the two systems where we demonstrate our methodology, a fermionic chain and a quantum many-body spin-1/2 chain. The fermionic chain represents a single-particle model, and it allows us to demonstrate our methodology as a generalization of Hamiltonian extraction with quasiparticle interference. Beyond its application on a conventional metal, it is worth noting that a single particle Hamiltonian learning could be performed as a quasiparticle description of a quantum magnet in terms of auxiliary spinons. The quantum many-body spin chain solved with tensor network methods represents the more accurate and challenging case, and is the case where methodology demonstrates its potential. In this case, we compute the exact quantum many-body ground and required expectation values using the full quantum many-body wavefunction using tensor networks
@@ III.2 Quantum spin 1/2 model with a spin-1 impurity
where |$\ket{\Omega({B_{z}})}$ is the ground state of the system at field $B_{z}$, $S^{z}_{i}$ is the z-component of the spin operator at site $n$, and $S^{z}_{\alpha}$ is the z-component of the spin operator at the impurity site $\alpha$. This observable is related to the response of the system to local Ising impurities as described in Sec. II, and in the gapped limit is given directly by the magnetization in the presence of a perturbation as shown in Eq. 6. We use the static correlator with respect to the impurity in the middle of the chain, where we considered chains with $L=21$ sites, with the central site the $S=1$ site.
IV Methods
In this section, we present the methodologies employed in our study to explore Hamiltonian learning using machine learning techniques.
IV.1 Hamiltonian Learning with neural networks
Our objective is to develop an algorithm capable of mapping the field-dependent non-local correlators to the Hamiltonian of the system. While the inverse step can be performed systematically with a quantum many-body solver, extracting the Hamiltonian from observables requires an alternative strategy. We use a machine learning methodology based on supervised learning for this purpose, for which a trainign set with Hamiltonians and non-local correlators is required. In Hamiltonian learning, the targets are specifically the parameters of the Hamiltonian that define the interactions within the system. The Hamiltonian we will study take the general form of
@@ IV.1 Hamiltonian Learning with neural networks
where $\mathcal{H}_{n}$ represents the interaction terms with their corresponding parameters $\Lambda_{n}$. This approach allows us to easily add or remove interactions from our model as needed. The targets, or the parameters $\Lambda_{n}$, are initially sampled uniformly within a specific range. This range is carefully chosen to avoid crossing any phase transitions, ensuring the stability and accuracy of our model. After feeding the NN with both input data and targets, the model should be trained and ready to make predictions on a brand new test set of data.
We start with the one-dimensional electron gas interacting sites with an impurity (Fig. 1b). The presence of scatterers allows us to probe the response of the system to perturbations, leading to frequency and spatial features that the NN leverages to extract the Hamiltonian. We consider first and second neighbour interactions between the sites and, most importantly, the spin-orbit coupling (SOC). SOC terms present significant challenges for NN learning due to their intrinsic complexity. These terms involve non-linear and anisotropic interactions between spin and orbital angular momentum, increasing the dimensionality of the problem. Additionally, SOC effects are highly dependent on the system’s symmetry, making generalization difficult for the NN. The introduction of SOC also leads to complex electron correlation effects, which are not easily captured by NNs.
@@
IV.2 Neural Network Architecture
[…]
IV.3 Fidelity Evaluation
To assess the robustness of the NN training against noise, we define fidelity using the Pearson correlation coefficient [34, 71, 72]:
@@ IV.4 Noise models
V Noisy Hamiltonian learning
@@
V.1 Single-particle system
@@
Figure 5: Evolution of the static spin correlator $\chi$ as a function of the external magnetic field $B^{z}$ (a,c) and the external magnetic field $B^{x}$ (b,d). Panels (a,b) show the pristine spin correlators, whereas panels (c,d) show the noisy correlators with magnetic noise at $\Delta_{B}^{2}=0.1$. The evolution of the correlator with the field acts as fingerprint of the many-body system, which the machine learning algorithm leverages to extract the Hamiltonian.
V.2 Hamiltonian learning a many-body quantum spin model
@@
Both directions of the magnetic field reflect a non-trivial evolution of the group state, and for our purpose any of those will enable us to perform Hamiltonian learning. In the configuration with $B_{x}$, the static spin correlator shows distinct correlation peaks centered around the middle impurity site, which are more symmetric and localized in the noiseless data. When noise is introduced, the clarity of these features decreases but remains discernible. Compared to the $B_{z}$ configuration, the $B_{x}$ field produces different correlation patterns, reflecting the anisotropic nature of the system. Both configurations demonstrate that, although noise impacts the sharpness of the data, key features near the impurity site are robust enough to be used for further analysis, such as training neural networks to infer Hamiltonian parameters.
@@
In Fig. 6b, the fidelity of the spin chain model parameters $J_{2}$, $J_{z}$, and $J_{DM}$ is presented as a function of the noise level $W$ for the case where the magnetic field $B$ is applied along the $z$-direction. This comparison includes three scenarios: perfect data,$\Delta_{B}^{2}=0.05$ and $\Delta_{B}^{2}=0.1$ magnetic noise level. As observed with the fermionic chain, the fidelity decreases with increasing noise levels. The robustness of $J_{2}$ and $J_{z}$ is evident, maintaining high fidelity even with $\Delta_{B}^{2}=0.1$ up to $W=2$. However, $J_{DM}$ shows greater sensitivity to noise, with fidelity dropping significantly as the noise increases, highlighting the stronger impact of noise on learning this parameter. Additionally, magnetic noise broadly diminishes fidelity across all parameters, underscoring its significant effect on model performance.
We show in Fig. 6c the results for the magnetic field $B$ applied along the $x$-direction. The overall trend remains similar, with fidelity decreasing as noise levels increase. However, the fidelity across parameters in the $B_{x}$ case shows slightly different sensitivities to noise compared to the $B_{z}$ case. This difference is likely correlated to the lack of variation in the datasets when $B$ is aligned along the $x$-axis, suggesting that the reduced variation in the input data could be influencing the model’s ability to learn and generalize across different noise levels. Finally, it is worth noting that more advanced machine learning methodologies such as GANs [86, 47] and diffusion models [87, 88] could further improve the efficiency of Hamiltonian learning.
VI Conclusion

### v2

# Hamiltonian-learning quantum magnets with non-local impurity tomography

@@ I Introduction
Quantum magnetism provides a highly flexible playground to engineer exotic phenomena [1, 2]. While a variety of natural materials host frustrated and magnetic states, accessing the most unconventional regimes often requires using artificial platform where Hamiltonians can be engineered [3]. Atomic-scale quantum magnets [4, 5, 6, 7, 8, 9, 10, 11] have opened new avenues for exploring fundamental quantum phenomena, enabling controllable platform to engineer symmetry breaking [12], quantum criticality [13], emergent excitations [14, 15], and topological states [16, 17]. Scanning tunneling microscopy (STM) enables creating and probing artificial quantum states with high spatial and energy resolution, providing detailed insights into the behavior of quantum magnets [4, 5, 6, 7, 8, 18, 19, 9, 10, 11]. Electrically driven paramagnetic resonance with STM has brought quantum control sensing at the atomic scale to remarkable levels. ESR-STM enabling energy resolution several orders of magnitude below thermal noise [10, 20], allowed quantum gate at the atomic level [21], and probing time evolution of magnetic dynamics with atomic accuracy [22]. Single atom manipulation in quantum magnets thus enables a degree of control radically different than in bulk materials, motivating the development of new strategies to tackle open problems in quantum magnetism [23].
@@
Understanding an experimental quantum many-body system relies on having an accurate description of the many-body interactions. Traditional methods of determining Hamiltonian parameters often rely on fitting experimental data to theoretical models. Characterizing quantum spin liquids represents a remarkable problem in quantum materials [1, 2] due to the intricate nature of their ground state, often requiring new strategies to find their experimental signatures [24, 25, 26, 27, 28, 29, 30]. As a result, even in relatively simple quantum magnets, determining the Hamiltonian from experimental data can be remarkably difficult [3]. Machine learning provides a flexible strategy to enable learning of Hamiltonian parameters directly from physical observables [23, 31, 32, 33, 34]. Hamiltonian learning can be done with data generated from theoretical models, that can be later on directly applied to experimental measurements [35, 36]. This strategy exploits the exceptional ability of machine learning methods to identify non-trivial correlations in observables that enable extracting information about the underlying ground state and Hamiltonian [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49].
Recent works have demonstrated that theoretical studies in Hamiltonian learning and spin systems can have direct implications for experimental investigations, particularly in scanning probe microscopy (SPM) and quantum materials characterization [50, 23, 51, 52, 53, 36]. For instance, theoretical advances in machine learning applied to Hamiltonian extraction have found applications in topological magnetic states [50], confined quantum magnets [23], molecular spin qubits [51], and even artificial nanoscale superconductors [36]. Our work follows this line of research, showing that machine learning techniques can be exploited to extract Hamiltonian parameters directly from local magnetization measurements. This approach not only aligns with recent efforts in theoretical Hamiltonian learning but also presents a concrete methodology that can be applied to experimental data.
@@
Here we show how the Hamiltonian of a quantum magnet can be directly inferred from measurements of the response of the spin chain to individual impurities. Our strategy relies on using the response of the system to extract the Hamiltonian using a machine learning strategy. The inclusion of individual impurities creates a perturbation in the ground state that directly reflects the underlying Hamiltonian, a correlation that can be mapped with a supervised neural network (NN). We show that this methodology enables extracting long-range and anisotropic exchange interactions, often a challenge to infer in experiments. We show that this methodology can be understood as a generalization of Hamiltonian extraction in a fermionic model, a procedure that is often performed in STM via quasiparticle interference. Our manuscript is organized as follows. In Section II, we describe the experimental strategy to obtain the data required for our Hamiltonian extraction. Section III details the many-body spin models we study, including the fermionic version that provides a minimal exemplification of our methodology. The methodology employed for Hamiltonian parameter extraction using NNs is described in Section IV. In Section V, we discuss the inclusion of magnetic and scaling noise in the datasets to simulate experimental conditions. In Section VI, we present the results of the Hamiltonian inference, demonstrating the resilience of our methodology to noise both in the evaluation as well as during training. Finally, in Section VII, we summarize our conclusions.
@@ II Impurity perturbations and non-local response
Our methodology will rely on using the non-local response of the system to extract the Hamiltonian. However, STM measures local quantities rather than non-local ones, and thus first an strategy must be developed to extract the required non-local information. Non-local correlators of quantum spin chain determine the response of the quantum many-body system to perturbation in the sites, and thus are reflected when adding additional impurities that act as exchange sources. Specifically, additional sites feature strong single-ion anisotropy as local perturbations, which then enable to measure elsewhere in the system what the change they created to the ground state. Experimentally, this was demonstrated with dysprosium single atom magnets [54], enabling atomic-scale magnetic field with stabilities above days. The Dy impurity creates a local exchange field in the neighboring site, which as a result modified the expectation value of the magnetization in the rest of the system. This strategy is exemplified in Fig. 1a.
@@
$H_{0}$ $=\sum_{n}J_{1}(1+\delta(-1)^{n})\boldsymbol{S}_{n}\cdot\boldsymbol{S}_{n}+$
$+\sum_{k=2}^{5}\sum_{n}J_{k}\boldsymbol{S}_{n}\cdot\boldsymbol{S}_{n+k},$ (7)
which features a many-body gap for $\delta\neq 0$ and $J_{k}\ll J_{1}$ for $k=2,3,4,5$. Figure 2a shows the magnetization in the different sites upon adding the local exchange perturbation, whereas Figure 2b shows the non-local spin-spin correlator in the absence of perturbation. It is clearly observed that both quantities feature very good correspondance over several orders of magnitude, which allows to infer non-local correlators from measuring local magnetizations of perturbed quantum magnet. The system features a decay correlation and magnetization for all the different couplings considered, yet interestingly the type of long range coupling included dramatically changes the speed of the decay and the oscillations. In this fashion, even though all those system feature similar many-body gaps, the magnetization and correlators provide a strong fingerprint of the underlying Hamiltonian. These unique behavior for the different coupling parameters is what provides the starting ground to attempt Hamiltonian learning from the observables in Figure 2.
The previous methodology exemplifies how non-local spin correlations can be extracted from local STM measurements by engineering local magnetic sites [4, 5, 6, 59]. As a result, Hamiltonian learning could be performed either directly by measuring the local response a system when adding local perturbations on different sites (Fig. 2a), or with the non-local correlator in the unperturbed limit (Fig. 2b). It is worth noting that this strategy can also be used for single particle systems, just by replacing the spin operators by fermionic operators and adapting the Kubo formalism accordingly. In terms of Hamiltonian learning, we will illustrate both single particle and the quantum many-body magnet cases. For the sake of concreteness, we will perform Hamiltonian learning with local measurements of perturbed systems for the fermionic case, and with the non-local correlators for the quantum many-body limit, yet noting that either local perturbed measurements or non-local unperturbed measurements allow for Hamiltonian learning in both cases. It is finally worth noting that time-dependent measurements, either local or non-local, could be used for Hamiltonian learning [60, 61, 62, 63, 64, 65, 66, 67, 68, 69].
@@ III Models
In this section, we introduce the two systems where we demonstrate our methodology, a fermionic chain and a quantum many-body spin-1/2 chain. The fermionic chain represents a single-particle model, and it allows us to demonstrate our methodology as a generalization of Hamiltonian extraction with quasiparticle interference. Beyond its application on a conventional metal, it is worth noting that a single particle Hamiltonian learning could be performed as a quasiparticle description of a quantum magnet in terms of auxiliary spinons. The quantum many-body spin chain solved with tensor network methods represents the more accurate and challenging case, and is the case where methodology demonstrates its potential. In this case, we compute the exact quantum many-body ground and required expectation values using the full quantum many-body wavefunction using tensor networks [70, 71].
@@ III.2 Quantum spin 1/2 model with a spin-1 impurity
where |$\ket{\Omega({B_{z}})}$ is the ground state of the system at field $B_{z}$, $S^{z}_{i}$ is the z-component of the spin operator at site $n$, and $S^{z}_{\alpha}$ is the z-component of the spin operator at the impurity site $\alpha$. This observable is related to the response of the system to local Ising impurities, as described in Sec. II. In the gapped limit, it is directly given by the magnetization in the presence of a perturbation, as shown in Eq. 6. Figure 2 presents results with long-range interactions, demonstrating that the relationship between non-local correlators and the system’s response to a perturbation holds in these generic models.
For Hamiltonian learning, we focus on the experimentally relevant scenario where $J_{1}$, $J_{2}$, $J_{DM}$, and $J_{z}$ are present, corresponding to the determination of four non-trivial parameters of the spin Hamiltonian.
We compute the static correlator with respect to the impurity placed at the center of the chain, considering systems of length $L=21$, where the central site corresponds to a $S=1$ spin.
IV Hamiltonian Learning with Neural Networks
In this section, we present the methodologies used in our study to explore Hamiltonian learning using machine learning techniques.
Our objective is to develop an algorithm capable of mapping the field-dependent non-local correlators to the Hamiltonian of the system [78, 79, 80]. While the inverse step can be performed systematically with a quantum many-body solver, extracting the Hamiltonian from observables requires an alternative strategy. We use a machine learning methodology based on supervised learning for this purpose, for which a training set with Hamiltonians and non-local correlators is required. In Hamiltonian learning, the targets are specifically the parameters of the Hamiltonian that define the interactions within the system.
The Hamiltonian we study follows the general form:
@@ IV Hamiltonian Learning with Neural Networks
where $\mathcal{H}_{n}$ represents the interaction terms with their corresponding parameters $\Lambda_{n}$. This formulation allows for flexibility in adding or removing interactions based on the specific physical system under consideration.
The target parameters $\Lambda_{n}$ are initially sampled uniformly within a specific range, chosen to ensure that the system remains within a stable phase and does not cross phase transitions to a symmetry broken regime. When selecting the dataset, it is essential to ensure that the measured spin correlations provide sufficient information for reliable parameter inference.
The choice of the Hamiltonian model primarily depends on the specific atom hosting the magnetic moment, the substrate used, and the spacing between atoms. As a concrete example, Ti atoms in MgO [20, 8, 81, 17] realize a model with finite $J_{1}$, a contribution from $J_{2}$, and very small contributions from $J_{Z}$ and $J_{\text{DM}}$. Both $J_{Z}$ and $J_{\text{DM}}$ stem from spin-orbit coupling effects, which are relatively weak in Ti and MgO. In contrast, replacing Ti with Ho [82] would significantly enhance the contributions of $J_{Z}$ and $J_{\text{DM}}$, due to the stronger spin-orbit coupling. More generally, for spins on a surface, the sign and magnitude of $J_{1}$ and $J_{2}$ can vary depending on the specific atomic arrangement [83]. This flexibility allows for tuning regimes where $J_{Z}$ and $J_{\text{DM}}$ are comparable to, or even stronger than, $J_{1}$ and $J_{2}$, by an appropriate choice of atom, substrate, and site arrangements [84].
In our case, we adopt a widely used Hamiltonian for spin systems, incorporating exchange interactions $J_{1},J_{2},J_{Z},J_{\text{DM}}$, which are relevant in many quantum magnetic materials, as described in Section III.2. A key advantage of our approach is its adaptability: the framework can be applied to various experimental setups by modifying the Hamiltonian terms to match the relevant physical interactions. To ensure experimental applicability, the training dataset must be carefully selected so that it covers a broad range of realistic parameters. Our model accounts for diverse experimental scenarios by incorporating different choices of atoms, substrates, and geometric arrangements, ensuring robustness across various conditions.
We start with the one-dimensional electron gas interacting sites with an impurity (Fig. 1b). The presence of scatterers allows us to probe the response of the system to perturbations, leading to frequency and spatial features that the NN leverages to extract the Hamiltonian. We consider first- and second-neighbor interactions between the sites and, most importantly, the spin-orbit coupling (SOC). SOC terms present significant challenges for NN learning due to their intrinsic complexity. These terms involve non-linear and anisotropic interactions between spin and orbital angular momentum, increasing the dimensionality of the problem. Additionally, SOC effects are highly dependent on the system’s symmetry, making generalization difficult for the NN. The introduction of SOC also leads to complex electron correlation effects, which are not easily captured by NNs.
@@
The details of the neural network training for both fermionic chain and spin chain can be found in Appendix A.
IV.1 Fidelity Evaluation
To assess the robustness of the NN training against noise, we define the fidelity using the Pearson correlation coefficient [34, 85, 86]:
@@ V Noise models
Figure 5: Evolution of the static spin correlator $\chi$ as a function of the external magnetic field $B^{z}$ (a,c) and the external magnetic field $B^{x}$ (b,d). Panels (a,b) show the pristine spin correlators, whereas panels (c,d) show the noisy correlators with magnetic noise at $\Delta_{B}^{2}=0.1$. The evolution of the correlator with the field acts as fingerprint of the many-body system, which the machine learning algorithm leverages to extract the Hamiltonian. The $n$-axis represents the site index of the spins in the chain.
VI Noisy Hamiltonian learning
@@
VI.1 Single-particle system
@@
VI.2 Hamiltonian learning a many-body quantum spin model
@@
Both directions of the magnetic field reflect a non-trivial evolution of the group state, and for our purpose any of those will enable us to perform Hamiltonian learning. In the configuration with $B_{x}$, the static spin correlator shows distinct correlation peaks centered around the middle impurity site, which are more symmetric and localized in the noiseless data. When noise is introduced, the clarity of these features decreases but remains discernible. Compared to the $B_{z}$ configuration, the $B_{x}$ field produces different correlation patterns, reflecting the anisotropic nature of the system. Both configurations demonstrate that, although noise impacts the sharpness of the data, key features near the impurity site are robust enough to be used for further analysis, such as training NNs to infer Hamiltonian parameters.
@@
In Fig. 6b, we present the fidelity of the spin chain model parameters $J_{2}$, $J_{z}$, and $J_{DM}$ as a function of the noise level $W$ when the magnetic field $B$ is applied along the $z$-direction. This analysis includes three scenarios: perfect data, $\Delta_{B}^{2}=0.05$, and $\Delta_{B}^{2}=0.1$ magnetic noise level. As observed with the fermionic chain, fidelity decreases with increasing noise. The parameters $J_{2}$ and $J_{z}$ exhibit robustness, maintaining high fidelity even for $\Delta_{B}^{2}=0.1$ up to $W=2$. Conversely, $J_{DM}$ is more sensitive, with a sharp fidelity drop at higher noise levels, highlighting its stronger susceptibility to perturbations. In general, magnetic noise further reduces fidelity across all parameters, confirming its substantial impact on learning accuracy.
Figure 6c shows results for $B$ applied along the $x$-direction. The overall trend remains similar, with fidelity decreasing as noise increases. However, the parameters display slightly different sensitivities compared to the $B_{z}$ case. This is likely due to the reduced variation in the dataset when $B$ is aligned along the $x$-axis, which may limit the model’s ability to extract relevant features under noisy conditions.
The effectiveness of Hamiltonian learning across different interaction regimes is further supported by our results, even in the presence of frustration and gapless spectra. The $J_{1}$-$J_{2}$ model, known to exhibit spin-liquid behavior in the antiferromagnetic regime [99, 100], poses challenges due to the potential absence of long-range order. However, our methodology remains reliable in these conditions, thanks to the inclusion of different frustration levels in the training data. This ensures that the model generalizes well across parameter regions, maintaining high fidelity despite minor variations in accuracy.
A crucial factor affecting the robustness of Hamiltonian learning to noise is the presence of a many-body gap. In gapped systems, low-energy excitations are suppressed, leading to short-range correlations that naturally filter out high-energy noise. This effect enhances the stability of inferred parameters, reducing the impact of fluctuations. In contrast, gapless systems exhibit long-wavelength modes, which introduce extended correlations and amplify noise effects. Consequently, parameter reconstruction may become less accurate, especially when the noise amplitude approaches or exceeds the system’s intrinsic energy scales. Nevertheless, our numerical results indicate that these gap-dependent noise effects do not significantly degrade the overall performance of our machine learning approach, which remains effective across the entire parameter space.
Finally, we note that more advanced machine learning techniques, such as GANs [101, 48] and diffusion models [102, 103], could further enhance the efficiency of Hamiltonian learning.
VII Conclusion